| #10 |
14 Jan 09:02 Experiment Configuration change
|
new
|
|
normal
|
2015_01 VIPIR + Topside
|
ISR
|
| #36 |
10 Feb 18:00 Imaging experiment is started
|
new
|
I. Manay
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR
|
| #49 |
11 Feb 11:11 Lost of data
|
new
|
A. Aguilar
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR
|
| #57 |
11 Feb 20:31 Problem with PC ARES
|
new
|
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR
|
| #58 |
11 Feb 20:32 Faraday Desincronization
|
new
|
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR
|
| #59 |
11 Feb 20:32 Restart EW_Drift
|
new
|
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR
|
| #60 |
11 Feb 20:40 Restart Faraday
|
new
|
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR
|
| #62 |
12 Feb 01:40 OFF Imaging
|
new
|
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR
|
| #72 |
13 Feb 00:30 Imaging system shut down, EW drift RF signal reconnected
|
new
|
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR+Imaging
|
| #82 |
14 Feb 01:30 Imaging experiment is finished
|
new
|
I. Manay
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR
|
| #87 |
14 Feb 11:50 Attenuation Channel E
|
new
|
AVALDEZ
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR
|
| #89 |
14 Feb 22:40 Restart Adquisition
|
new
|
P.Condor
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR
|
| #93 |
14 Feb 19:50 Faraday Desynchronization
|
new
|
P. Condor
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR
|
| #96 |
14 Feb 22:00 The problem was the pulse of EW_Drifts
|
new
|
P. Condor
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR
|
| #97 |
14 Feb 22:40 Restart Acquisition
|
new
|
P. Condor
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR
|
| #99 |
15 Feb 02:07 Imaging OFF
|
new
|
P. Condor
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR
|
| #102 |
15 Feb 06:40 Faraday Desynchronization
|
new
|
P. Condor
|
normal
|
2015_02 ISR + Imaging (ISCOD)
|
ISR
|
| #133 |
25 Mar 01:00 Imaging Off
|
new
|
|
normal
|
2015_03 ISR + Imaging CNOFS Support
|
ISR
|
| #140 |
25 Mar 10:20 Faraday experiment unsynchronized
|
new
|
|
normal
|
2015_03 ISR + Imaging CNOFS Support
|
ISR+Imaging
|
| #141 |
25 Mar 10:23 Ares PC rebooted, it was halted.
|
new
|
|
normal
|
2015_03 ISR + Imaging CNOFS Support
|
ISR+Imaging
|
| #144 |
25 Mar 10:50 Restarting clock divider device.
|
new
|
|
normal
|
2015_03 ISR + Imaging CNOFS Support
|
ISR+Imaging
|
| #145 |
25 Mar 10:23 Changing the Trendnet device that change the aquisition computers monitors.
|
new
|
|
normal
|
2015_03 ISR + Imaging CNOFS Support
|
ISR+Imaging
|
| #146 |
25 Mar 11:20 System completely restored with the processing software.
|
new
|
|
normal
|
2015_03 ISR + Imaging CNOFS Support
|
ISR+Imaging
|
| #148 |
25 Mar 17:40 Imaging system started.
|
new
|
|
normal
|
2015_03 ISR + Imaging CNOFS Support
|
ISR+Imaging
|
| #153 |
26 Mar 01:00 Imaging experiment ended
|
new
|
I. Manay
|
normal
|
2015_03 ISR + Imaging CNOFS Support
|
ISR
|
| #166 |
27 Mar 00:45 Imaging experiment ended
|
new
|
|
normal
|
2015_03 ISR + Imaging CNOFS Support
|
ISR+Imaging
|
| #175 |
26 Mar 22:00 Imaging Acquisition System Lost of synchronization
|
new
|
ryanque
|
normal
|
2015_03 ISR + Imaging CNOFS Support
|
ISR
|
| #203 |
14 Abr 20:00 Ch B had high noise
|
new
|
P. Condor
|
normal
|
2015_04_1 MST - ISR - EEJ
|
ISR
|
| #213 |
15 Abr 13:40 Acquisition software halted
|
new
|
|
normal
|
2015_04_1 MST - ISR - EEJ
|
ISR
|
| #214 |
15 Abr 14:00 Acquisition software halted
|
new
|
|
normal
|
2015_04_1 MST - ISR - EEJ
|
ISR
|
| #216 |
15 Abr 16:00 Acquisition system hard drive changed.
|
new
|
|
normal
|
2015_04_1 MST - ISR - EEJ
|
ISR
|
| #217 |
15 Abr 16:10 End of the hard drive change.
|
new
|
|
normal
|
2015_04_1 MST - ISR - EEJ
|
ISR
|
| #257 |
26 May 17:49 EW Drift Acquisition stops
|
new
|
|
normal
|
2015_05 ISR + Imaging
|
ISR
|
| #262 |
27 May 00:10 Imaging experiment stopped by us.
|
new
|
|
normal
|
2015_05 ISR + Imaging
|
ISR+Imaging
|
| #275 |
28 May 00:40 Imaging experiment stopped by us.
|
new
|
|
normal
|
2015_05 ISR + Imaging
|
ISR
|
| #284 |
29 May 00:36 Imaging Off
|
new
|
P. Condor
|
normal
|
2015_05 ISR + Imaging
|
ISR
|
| #296 |
09 Jun 01:21 We change TR of Ch E
|
new
|
P. Condor
|
normal
|
2015_06 Valley Experiment
|
ISR
|
| #297 |
26 May 00:38 Reset Pc Ares
|
new
|
P. Condor
|
normal
|
2015_05 ISR + Imaging
|
ISR
|
| #300 |
25 May 23:26 Change of cable ChB
|
new
|
P. Condor
|
normal
|
2015_05 ISR + Imaging
|
ISR
|
| #301 |
25 May 22:20 Test in Channel B and D -EW_Drifts
|
new
|
P. Condor
|
normal
|
2015_05 ISR + Imaging
|
ISR
|
| #302 |
25 May 22:15 Change of TR -Faraday
|
new
|
P. Condor
|
normal
|
2015_05 ISR + Imaging
|
ISR
|
| #306 |
09 Jun 07:48 Attenuation ch E
|
new
|
P. Condor
|
normal
|
2015_06 Valley Experiment
|
ISR
|
| #307 |
10 Jun 08:10 Restart Adquisition
|
new
|
P. Condor
|
normal
|
2015_06 Valley Experiment
|
ISR
|
| #321 |
12 Jun 07:36 Plots coherence and cross spectra
|
new
|
P. Condor
|
normal
|
2015_06 Valley Experiment
|
ISR
|
| #336 |
08 Jul 08:30 Reset JARS System
|
new
|
|
normal
|
2015_07 MST-ISR-EEJ
|
ISR
|
| #339 |
08 Jul 13:58 change drive system PC-JARS
|
new
|
|
normal
|
2015_07 MST-ISR-EEJ
|
ISR
|
| #351 |
10 Jul 00:32 Acquisition system was re-started
|
new
|
I. Manay
|
normal
|
2015_07 MST-ISR-EEJ
|
ISR
|
| #352 |
10 Jul 06:34 Acquisition was restarted
|
new
|
I. Manay
|
normal
|
2015_07 MST-ISR-EEJ
|
ISR
|
| #408 |
20 Jul 16:30 Start Imaging
|
new
|
P. Condor
|
normal
|
2015_07_2 Support ISR School
|
ISR
|
| #409 |
21 Jul 08:00 Imaging Off and restart Faraday
|
new
|
P. Condor
|
normal
|
2015_07_2 Support ISR School
|
ISR
|
| #410 |
21 Jul 08:08 Ch B - Faraday is attenuated 2 db
|
new
|
P. Condor
|
normal
|
2015_07_2 Support ISR School
|
ISR
|
| #411 |
21 Jul 08:09 We changed ChA for Ch B faraday
|
new
|
P. Condor
|
normal
|
2015_07_2 Support ISR School
|
ISR
|
| #428 |
22 Jul 08:20 Restart Faraday
|
new
|
P. Condor
|
normal
|
2015_07_2 Support ISR School
|
ISR
|
| #429 |
22 Jul 10:00 Reset batch Faraday for desinchronization
|
new
|
P. Condor
|
normal
|
2015_07_2 Support ISR School
|
ISR
|
| #430 |
22 Jul 11:00 Attenuation in Faraday
|
new
|
P. Condor
|
normal
|
2015_07_2 Support ISR School
|
ISR
|
| #431 |
22 Jul 14:59 Reset pc Faraday
|
new
|
P. Condor
|
normal
|
2015_07_2 Support ISR School
|
ISR
|
| #432 |
22 Jul 16:15 Restart Faraday
|
new
|
P. Condor
|
normal
|
2015_07_2 Support ISR School
|
ISR
|
| #433 |
22 Jul 17:31 Start Imaging
|
new
|
P. Condor
|
normal
|
2015_07_2 Support ISR School
|
ISR
|
| #440 |
23 Jul 06:30 Stop Imaging
|
new
|
|
normal
|
2015_07_2 Support ISR School
|
ISR
|
| #467 |
26 Ago 20:26 Faraday loss synchronization
|
new
|
P. Condor
|
normal
|
2015_08 Support CNOFS
|
ISR
|
| #476 |
27 Ago 07:10 Turn off Imaging
|
new
|
P. Condor
|
normal
|
2015_08 Support CNOFS
|
ISR
|
| #477 |
27 Ago 07:20 Noise levels
|
new
|
P. Condor
|
normal
|
2015_08 Support CNOFS
|
ISR
|
| #481 |
27 Aug 08:16 Faraday acquisition was restarted
|
new
|
I. Manay
|
normal
|
2015_08 Support CNOFS
|
ISR+Imaging
|
| #482 |
27 Aug 11:44 Faraday acquisition was restarted
|
new
|
I. Manay
|
normal
|
2015_08 Support CNOFS
|
ISR+Imaging
|
| #499 |
22 Sep 11:17 Faraday acquisition was restarted
|
new
|
I. Manay
|
normal
|
2015_09_1 Support CNOFS
|
ISR+Imaging
|
| #506 |
23 Sep 00:06 Imaging Off
|
new
|
|
normal
|
2015_03 ISR + Imaging CNOFS Support
|
ISR+Imaging
|
| #522 |
23 Sep 18:00 Imaging system turned on
|
new
|
|
normal
|
2015_09_1 Support CNOFS
|
ISR+Imaging
|
| #527 |
23 Sep 21:50 Restart Faraday Pc
|
new
|
P. Condor
|
normal
|
2015_09_1 Support CNOFS
|
ISR
|
| #528 |
23 Sep 22:10 Restart Faraday
|
new
|
P. Condor
|
normal
|
2015_09_1 Support CNOFS
|
ISR
|
| #529 |
23 Sep 22:14 Reset of Faraday Pc
|
new
|
P. Condor
|
normal
|
2015_09_1 Support CNOFS
|
ISR
|
| #530 |
23 Sep 23:15 Restart EW Drift and Faraday batches
|
new
|
P. Condor
|
normal
|
2015_09_1 Support CNOFS
|
ISR
|
| #533 |
24 Sep 01:43 Off Imaging
|
new
|
P. Condor
|
normal
|
2015_09_1 Support CNOFS
|
ISR
|
| #535 |
24 Sep 05:50 Restart EW Drifts and faraday batches
|
new
|
P. Condor
|
normal
|
2015_09_1 Support CNOFS
|
ISR
|
| #605 |
10 Dec 09:30 Off DDS Imaging
|
new
|
P. Condor
|
normal
|
2015_12 ISR ISCOD
|
ISR
|
| #606 |
10 Dec 12:27 Off Transmitter
|
new
|
P. Condor
|
normal
|
2015_12 ISR ISCOD
|
ISR
|
| #609 |
10 Dec 14:19 On VIPIR with new IPP
|
new
|
P. Condor
|
normal
|
2015_12 ISR ISCOD
|
ISR
|
| #611 |
10 Dec 18:20 Attenuated ch 5 imaging
|
new
|
P. Condor
|
normal
|
2015_12 ISR ISCOD
|
ISR
|
| #643 |
13 Dec 03:11 OFF IMAGING
|
new
|
p
|
normal
|
2015_12 ISR ISCOD
|
ISR
|
| #656 |
09 Dec 19:22 Interference observed at the B channel of EW_Drift experiment
|
new
|
A. Valdez
|
normal
|
2015_12 ISR ISCOD
|
ISR+Imaging
|
| #676 |
05 Ene 17:36 Imaging On
|
new
|
P Condor
|
normal
|
2016_01_1 ISR+Imaging (ISCOD)
|
ISR
|
| #678 |
05 Ene 17:49 Tx on
|
new
|
P Condor
|
normal
|
2016_01_1 ISR+Imaging (ISCOD)
|
ISR
|
| #700 |
08 Jan 01:00 Turn Off Imaging
|
new
|
AValdez
|
normal
|
2016_01_1 ISR+Imaging (ISCOD)
|
ISR+Imaging
|
| #703 |
08 Jan 07:50 Operation Resume
|
new
|
AValdez
|
normal
|
2016_01_1 ISR+Imaging (ISCOD)
|
ISR+Imaging
|
| #710 |
09 Ene 02:10 Imaging Off
|
new
|
P.Condor
|
normal
|
2016_01_1 ISR+Imaging (ISCOD)
|
ISR
|
| #765 |
21 Ene 20:20 Restart batch in Jars
|
new
|
P.Condor
|
normal
|
2016_01_3 MST-ISR-EEJ
|
ISR
|
| #770 |
22 Ene 00:44 Reset Pc operaciones
|
new
|
P.Condor
|
normal
|
2016_01_3 MST-ISR-EEJ
|
ISR
|
| #792 |
22 Ene 17:00 JASMET processing begins
|
new
|
|
normal
|
2016_01_3 MST-ISR-EEJ
|
JASMET
|
| #793 |
22 Ene 18:47 JASMET processing fixed, the header of the processed data has CodeType=17, this seems to be the problem
|
new
|
|
normal
|
2016_01_3 MST-ISR-EEJ
|
JASMET
|
| #797 |
23 Jan 05:10 Change data Path
|
new
|
AVALDEZ
|
normal
|
2016_01_3 MST-ISR-EEJ
|
ISR
|
| #812 |
24 JAN 12:10 Problems with JARS
|
new
|
AValdez
|
normal
|
2016_01_3 MST-ISR-EEJ
|
ISR
|
| #817 |
25 Jan 07:00 Acq. re-starts
|
new
|
I. Manay
|
normal
|
2016_01_3 MST-ISR-EEJ
|
ISR
|
| #823 |
25 Jan 10:25 Acquisition system stopped
|
new
|
|
normal
|
2016_01_3 MST-ISR-EEJ
|
ISR
|
| #828 |
25 Jan 17:41 Acquisition restarted
|
new
|
|
normal
|
2016_01_3 MST-ISR-EEJ
|
ISR
|
| #831 |
25 Jan 19:18 Acquisition restarted
|
new
|
|
normal
|
2016_01_3 MST-ISR-EEJ
|
ISR
|
| #835 |
25 Jan 20:50 Acquisition restarted
|
new
|
|
normal
|
2016_01_3 MST-ISR-EEJ
|
ISR
|
| #836 |
25 Jan 22:14 Acquisition restarted
|
new
|
|
normal
|
2016_01_3 MST-ISR-EEJ
|
ISR
|
| #837 |
25 Jan 23:51 Plots
|
new
|
|
normal
|
2016_01_3 MST-ISR-EEJ
|
ISR
|
| #838 |
26 Jan 00:12 Acquisition restarted
|
new
|
|
normal
|
2016_01_3 MST-ISR-EEJ
|
ISR
|
| #839 |
26 Jan 00:21 Acquisition restarted
|
new
|
|
normal
|
2016_01_3 MST-ISR-EEJ
|
ISR
|
| #840 |
26 Jan 00:30 Acquisition restarted
|
new
|
|
normal
|
2016_01_3 MST-ISR-EEJ
|
ISR
|